Economics

Question 1:

FC VC TC AFC AVC ATC MC 920 475 60 What is the FC equal to? [Type a whole number, no gaps.]Question 2:

FC VC TC AFC AVC ATC MC 920 475 60 10 What is the MC at Q=2 equal to? [Type a whole number, no gaps.]
Answer #1

At Q = 2

TC = ATC *2 = 2*475 = $950.

So MC at Q = 2 is MC = $950 - $920 = $30 (Answer 2)

Now, we know that (FC + 3VC) = $950 and  (FC + 2VC) = $920.

Solving the above two equations for VC we get,

(920 - FC)/2 = (950 - FC)/3

or, 2760 - 3FC = 1900 - 2FC

or, FC = $860 (Answer 1)

Kindly provide a thumbs up for the answers.